Find three consecutive even integers so that 7 times the third is equal to 20 less than the first.
Answer by darq(90)

Find three consecutive even integers so that 7 times the third is equal to 20 less than the first.
x is the smallest
x + 2 is the medium
x + 4 is the greatest
7 (x + 4) = x - 20
7x + 28 = x - 20
6x = -48
x = -8
#1=-8
#2=-6
#3=-4